Operating Systems and Systems Integration

Laboratory Test

1 Scope

• This test covers all the concepts and skills taught in the laboratory classes concerning Samba

• Topics covered:

∘ Samba as a pdc with Windows 2000 as a member of the domain

∘ Creating users in the domain

∘ Using a common wins server

∘ Profiles for members of the domain

∘ Creating shares on the Samba server

∘ Creating useful log in scripts

∘ Use the Server Manager to manage the Samba server from the Windows 2000 domain member

∘ implementing Microsoft dfs (Distributed File System) on the Samba server

• Note that this is an open book test; you can refer to any material you like during the test.

2 What to do

• Form a team of two students. One student must have on their hard disk an installation of Windows 2000 that is not a primary or backup domain controller. It does not matter whether this is Windows 2000 (Professional or otherwise) or Windows 2000 Server (Advanced or otherwise).

• If at least one in your team does not have Windows 2000 as decribed above, then either install it, or choose another partner.

• Each team configures their two computers to meet the following requirements.

3 Context

A high school has a small budget with which to implement a reliable file and printing server for students. Each student has a home directory when they log in, and there are a number of shares to hold particular applications.

• The data should be accessible both from Windows 2000 and from Linux.

Laboratory Test Operating Systems and Systems Integration

2

• In the future there will be a Macintosh machine that will also need to access this data.

• Since reliability, cross-platform support and low cost are important criteria, and a number of staff are familiar with it, Linux was chosen as the platform.

• Create domain logon accounts for at least two students and one staff administrator, with logon scripts that map one common share as well as their own home directory to drive letters on the Windows clients.

• Create a group students and add the students to the group.

• Create a group user policy for students that:

∘ Hides the background tab to display properties, to prevent students from chang-

ing the desktop background ∘ Hides the settings tab to prevent students from changing the resolution ∘ Hides the Screen Saver tab to prevent students from changing the screen saver

settings ∘ Removes folders and taskbar from Settings on the Start Menu ∘ Remove the ShutDown Command from the Start Menu ∘ Don’t save settings at exit. ∘ Disable registry editing tools. ∘ Disable Lock Workstation ∘ Exclude Temporary Internet Files and Temp from the User’s Roaming Profile.

• Ensure that the domain Administrator account has no policy restrictions that reduce their administration rights.

• Create two dfs links on the Samba server: one that points to a share from the Samba server, and one that points to a share from the Windows machine that is a member of the Samba domain.

4 Assessment

• A group mark will be assigned to both team members, unless there is a dispute between members.

• You are expected to have completed the configuration before your lecturer comes to assess your work. Your lecturer may ask you to demonstrate how to configure particular items.

• A grade of A will be assigned to those who succefully complete most of the requirements, including the web server configuration.

• A grade of B will be assigned to those who complete most of the Samba configuration.

• A grade of fail will be assigned to those groups who are not able to set up the primary domain controller, or who are unable to show how to configure a number of settings.